OpenStudy (anonymous):

solve the system of equations* -2x-3y=14 -4x+2y=-4

OpenStudy (anonymous):

multiply the first one straight across by -2 then add to the second one

OpenStudy (anonymous):

need help fast takin a math test pleaseeeee

OpenStudy (anonymous):

shouldn't have said that.

OpenStudy (anonymous):

you get \[4x+6y=-24\]\[-4x+2y=-4\] add and get \[8y=-28\]

OpenStudy (anonymous):

oops i made a mistake

OpenStudy (anonymous):

\[4x+6y=-28\]\[-4x+2y=-4\] add and get\[8y=-32\] so \(y=-4\)

OpenStudy (anonymous):

if you just want the answer quick, don't ask here. you will get mistakes like the one i made just do this http://www.wolframalpha.com/input/?i=-2x-3y%3D14%2C+-4x%2B2y%3D-4